OpenStudy (anonymous):

Which expression is equivalent to (1 + 9x)10 ? A. –10 – 90x B. 11 + 9x C. 10 + 90x D. –11 – 9x

OpenStudy (anonymous):

Distribute 10 to1 and then to 9x..

OpenStudy (anonymous):

(1+9x)10= 10+90x....sistributive property.

OpenStudy (anonymous):

\[(1+9x) \cdot 10 = 1 \cdot 10 + 9x \cdot 10 = 10 + 90x\]